OpenStudy (anonymous):

Find the equation of the line that goes through the points (-15, 70) and (5,10)

OpenStudy (anonymous):

Solve for slope first using: y2-y1/x2-x1 to get: 70-10/5 -(-15) = 60/20 = 3 Plug that into: y = mx + b. Now, you have y = 3x + b. Use any of the ordered pairs you have to substitute into the equation. For instance, you use, (5,10) where x = 5 and y = 10 to get: 10 = 3(5) + b = 0 10 = 15 + b -5 = b *Note: subtract 15 from both sides So, now you have: y = 3x - 5